I would greatly appreciate it.1. Simplify each of the following Boolean functions using K-map a. F(a,b,c)={m(0,1,4,5,6,7) c. F(a,b,c)={m(0,1,2,4,5,6) i. F(
Answer #1

a fraib, c) Eloin 4,5,6,7) labo.0111110 : f=a+b © FlQjb,c) - Emloil, 2, 4,5,6) canal on the c! Facib (7) F(a,b,c)= S(O2, 44,

(m) f(a,b,c,d)=zm(6,8,10) colabodo u n Sbcd F = bcdtabd 100 ml sald (FV =ŽM1011, 23,6,7,10,11). - Dalby doof all 10 Oilu

w) Farbiced) = ECO,16A,4,13,2,10) f=cdthcltbld vo) Fanbio Carole 2(1,43,7) colabo, el 10 f=dbcd tabcolt dócd+abed